# Inference for multiple treatment effects using confounder importance learning

### v4

@@ E.1. Derivation of Equation (14)
which is independent of ${\bm{\theta}}$, and where $\pi_{t}$ is the marginal prior inclusion probability within our framework. Then, implementing the approximation (19) into (18) gives
(20) ${\bm{\theta}}^{\textsc{ep}}:=\operatorname*{arg\,max}_{{\bm{\theta}}\in\mathbb{R}^{T+1}}\sum_{({\bm{\gamma}},{\bm{\delta}})}h({\bm{\delta}})\times\\ \prod_{j=1}^{J}\left[r_{j}^{\textsc{ep}}\pi_{j}({\bm{\theta}})\right]^{\gamma_{j}}\left[(1-r_{j}^{\textsc{ep}})(1-\pi_{j}({\bm{\theta}}))\right]^{1-\gamma_{j}}.$
Note that the product in the RHS of (20) defines a probability distribution on $(\delta_{1},\dots,\delta_{T},\gamma_{1},\dots,\gamma_{J})$ with independent components, hence the sum is the normalizing constant of such distribution. Thus, this constant is just the product of the univariate normalizing constants. The univariate normalizing constant of each Bernouilli is then
@@
for every $r_{j}$, and similarly $s_{t}^{\textsc{ep}}\pi_{t}+(1-s_{t}^{\textsc{ep}})(1-\pi_{t})$ for every $s_{t}$. Hence, replacing into (20) we directly obtain
@@
E.2. Gradient of Equation (14)
@@
$\nabla_{{\bm{\theta}}}\sum_{j=1}^{J}\log h_{j}({\bm{\theta}})=(1-2\rho)\sum_{j=1}^{J}\mathbf{f}_{j}\frac{\pi_{j}({\bm{\theta}})(1-\pi_{j}({\bm{\theta}}))}{h_{j}({\bm{\theta}})}.$
@@
Appendix F Product MOM non-local prior
@@
Appendix G Computational Methods
G.1. Numerical computation of the marginal likelihood for non-local priors
@@
G.2. Numerical optimization for empirical Bayes hyper-parameters
Algorithm 1 describes our method to estimate ${\bm{\theta}}^{\textsc{ep}}$ and ${\bm{\theta}}^{\textsc{eb}}$. We employ the quasi-Newton BFGS algorithm to optimize the objective function. For ${\bm{\theta}}^{\textsc{eb}}$, we use the gradients from Proposition 3.1, while the Hessian is evaluated numerically using line search, with the R function nlminb. Note, however, that obtaining ${\bm{\theta}}^{\textsc{eb}}$ requires sampling models from their posterior distribution for each ${\bm{\theta}}$, which is impractical, to then obtain posterior inclusion probabilities required by (9). Instead, we restrict attention to the models $M$ sampled for either ${\bm{\theta}}=\mathbf{0}$ or ${\bm{\theta}}={\bm{\theta}}^{\textsc{ep}}$ in order to avoid successive MCMC runs at every step, relying on the relative regional proximity between the starting point ${\bm{\theta}}^{\textsc{ep}}$ and ${\bm{\theta}}^{\textsc{eb}}$. This proximity would ensure that $M$ contains the large majority of models with non-negligible posterior probability under ${\bm{\theta}}^{\textsc{eb}}$. For ${\bm{\theta}}^{\textsc{ep}}$, we use employ the same BFGS strategy using gradient computed in E.2, with numerical evaluation of the Hessian. This computation requires only one MCMC run at ${\bm{\theta}}=\mathbf{0}$, which allows us to use grid search to avoid local optima. As for the size of the grid, we let the user specify what points are evaluated. For $K$ points in the grid one must evaluate the log objective function $K^{T+1}$ times, so we recommend to reduce the grid density as $T$ grows. By default, we evaluate every integer in the grid assuming $T$ is not large, but preferably we avoid coordinates greater than 10 in absolute value, as in our experiments it is very unlikely that any global posterior mode far from zero is isolated, i.e. not reachable by BFGS by starting to its closest point in the grid. Additionally, even if that were the case, numerically it makes no practical difference, considering that marginal inclusion probabilities are bounded away from zero and one regardless.
Output: Estimates for ${\bm{\theta}}^{\textsc{ep}}$ and ${\bm{\theta}}^{\textsc{eb}}$ 1: Obtain $B$ posterior samples $({\bm{\gamma}},{\bm{\delta}})^{(b)}\sim p({\bm{\gamma}},{\bm{\delta}}\mid{\mathbf{y}},{\bm{\theta}}=\mathbf{0})$ for $b=1,\ldots,B$. Denote by $M^{(0)}$ the corresponding set of unique models.
@@ G.2. Numerical optimization for empirical Bayes hyper-parameters
Appendix H Supplementary Results
H.1. Illustration of the EB and EP objective functions

### v6

@@ 11. Derivations
which is independent of ${\bm{\theta}}$, and where $\pi_{t}$ is the marginal prior inclusion probability for treatment $t$ (by default, $\pi_{t}=1/2$). Then, taking (21) and replacing $p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})$ by the approximation given by (22) gives
(23) ${\bm{\theta}}^{\textsc{ep}}:=\operatorname*{arg\,max}_{{\bm{\theta}}\in\mathbb{R}^{T+1}}\sum_{({\bm{\gamma}},{\bm{\delta}})}h({\bm{\delta}})\prod_{j=1}^{J}\left[r_{j}^{\textsc{ep}}\pi_{j}({\bm{\theta}})\right]^{\gamma_{j}}\left[(1-r_{j}^{\textsc{ep}})(1-\pi_{j}({\bm{\theta}}))\right]^{1-\gamma_{j}}.$
The terms inside the sum in the right-hand side defines a probability distribution on $(\delta_{1},\dots,\delta_{T},\gamma_{1},\dots,\gamma_{J})$ with independent Bernoulli components, hence their sum is the normalizing constant of said distribution. Since the distribution has independent components, the normalizing constant is just the product of the univariate normalizing constants. The univariate normalizing constant of each Bernouilli is then
@@
for every $r_{j}$, and similarly $s_{t}^{\textsc{ep}}\pi_{t}+(1-s_{t}^{\textsc{ep}})(1-\pi_{t})$ for every $s_{t}$. Hence, we directly obtain
@@
11.2. Gradient of Equation 3.9
Note first that, since $\pi_{j}({\bm{\theta}})$ is constant when $\tilde{\pi}_{j}({\bm{\theta}})\not\in(\underline{\rho},\bar{\rho})$, the gradient for such terms is zero. We hence focus on $j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})$, since in that case $\pi_{j}({\bm{\theta}})=\tilde{\pi}_{j}({\bm{\theta}})$.
@@
(25) $\nabla_{{\bm{\theta}}}\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\log h_{j}({\bm{\theta}})=\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\frac{\pi_{j}({\bm{\theta}})(1-\pi_{j}({\bm{\theta}}))}{h_{j}({\bm{\theta}})}.$
Finally, note that
$\pi_{j}({\bm{\theta}})(1-\pi_{j}({\bm{\theta}}))=\pi_{j}({\bm{\theta}})(r_{j}-r_{j}\pi_{j}({\bm{\theta}})+(1-r_{j})-\pi_{j}({\bm{\theta}})+r_{j}\pi_{j}({\bm{\theta}}))=\pi_{j}({\bm{\theta}})r_{j}-\pi_{j}({\bm{\theta}})h_{j}({\bm{\theta}}),$
giving that (25) is
$\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[\frac{\pi_{j}({\bm{\theta}})r_{j}}{h_{j}({\bm{\theta}})}-\pi_{j}({\bm{\theta}})\right]=\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[P^{EP}(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})\right].$
where
$P^{EP}(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})=\frac{\pi_{j}({\bm{\theta}})r_{j}}{h_{j}({\bm{\theta}})},$
@@
12. Product MOM non-local prior
@@
13. Computational Methods
13.1. Numerical computation of the marginal likelihood for non-local priors
@@
13.2. Numerical optimization for empirical Bayes hyper-parameters
Algorithm 1 describes our method to estimate ${\bm{\theta}}^{\textsc{ep}}$ and ${\bm{\theta}}^{\textsc{eb}}$. We employ the quasi-Newton BFGS algorithm to optimize the objective function. For ${\bm{\theta}}^{\textsc{eb}}$, we use the gradients from Proposition 3.1, while the Hessian is evaluated numerically using line search, with the R function nlminb. Note, however, that obtaining ${\bm{\theta}}^{\textsc{eb}}$ requires sampling models from their posterior distribution for each ${\bm{\theta}}$, which is impractical, to then obtain posterior inclusion probabilities required by (10). Instead, we restrict attention to the models $M$ sampled for either ${\bm{\theta}}=\mathbf{0}$ or ${\bm{\theta}}={\bm{\theta}}^{\textsc{ep}}$ in order to avoid successive MCMC runs at every step, relying on the relative regional proximity between the starting point ${\bm{\theta}}^{\textsc{ep}}$ and ${\bm{\theta}}^{\textsc{eb}}$. This proximity would ensure that $M$ contains the large majority of models with non-negligible posterior probability under ${\bm{\theta}}^{\textsc{eb}}$. For ${\bm{\theta}}^{\textsc{ep}}$, we use employ the same BFGS strategy using gradient computed in 11.2, with numerical evaluation of the Hessian. This computation requires only one MCMC run at ${\bm{\theta}}=\mathbf{0}$, which allows us to use grid search to avoid local optima. As for the size of the grid, we let the user specify what points are evaluated. For $K$ points in the grid one must evaluate the log objective function $K^{T+1}$ times, so we recommend to reduce the grid density as $T$ grows. By default, we evaluate every integer in the grid assuming $T$ is not large, but preferably we avoid coordinates greater than 10 in absolute value, as in our experiments it is very unlikely that any global posterior mode far from zero is isolated, i.e. not reachable by BFGS by starting to its closest point in the grid. Additionally, even if that were the case, numerically it makes no practical difference, considering that marginal inclusion probabilities are bounded away from zero and one regardless.
Output: Estimates for ${\bm{\theta}}^{\textsc{ep}}$ and ${\bm{\theta}}^{\textsc{eb}}$
1: Obtain $B$ posterior samples $({\bm{\gamma}},{\bm{\delta}})^{(b)}\sim p({\bm{\gamma}},{\bm{\delta}}\mid{\mathbf{y}},{\bm{\theta}}=\mathbf{0})$ for $b=1,\ldots,B$. Denote by $M^{(0)}$ the corresponding set of unique models.
@@ 13.2. Numerical optimization for empirical Bayes hyper-parameters
14. Derivation of Bayes factor asymptotics
Proposition 1(i) in Rossell and Telesca (2017) gives that the Bayes factor under the pMOM prior $p(\alpha_{t}\mid\delta_{t}=1,\phi)=\frac{\alpha_{t}^{2}}{\phi\tau/v_{t}}\text{N}(\alpha_{t};0,\phi\tau/v_{t})$ and $p(\beta_{j}\mid\gamma_{j}=1,\phi)=\frac{\beta_{j}^{2}}{\phi\tau/v_{j}}\text{N}(\beta_{j};0,\phi\tau/v_{j})$ is equal to
(26) $\frac{p({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}})}{p({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*})}=\frac{E^{N}\left(\prod_{\delta_{t}=1}\alpha_{t}^{2}v_{t}/[\phi\tau]\prod_{\gamma_{j}=1}\beta_{j}^{2}v_{j}/[\phi\tau]\mid{\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\right)}{E^{N}\left(\prod_{\delta_{t}^{*}=1}\alpha_{t}^{2}v_{t}/[\phi\tau]\prod_{\gamma_{j}^{*}=1}\beta_{j}^{2}v_{j}/[\phi\tau]\mid{\mathbf{y}},{\bm{\delta}}^{*},{\bm{\gamma}}^{*}\right)}\frac{p^{N}({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}})}{p^{N}({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}})}$
where
[…]
is the marginal likelihood under the Normal prior featuring in the pMOM density, $E^{N}()$ denotes a posterior expectation under said Normal prior, $V_{\delta}$ is diagonal with entries given by the $v_{t}$’s and $V_{\gamma}$ is diagonal with entries given by the $v_{j}$’s.
Asymptotics for (26) are obtained by studying separately the first term involving the ratio of posterior means, and the second term featuring the Bayes factor obtained under the Normal prior. Before presenting the arguments, we outline further notation and assumptions needed for the results to hold.
In this section, for simplicity we denote by $\bm{\zeta}=({\bm{\alpha}},{\bm{\beta}})$ the whole set of parameters and by $\bm{\zeta}_{\delta,\gamma}$ the subset of non-zero values under model $({\bm{\delta}},{\bm{\gamma}})$.
Assume that $y_{1},\ldots,y_{n}$ arise independently from an unknown data-generating distribution $F^{*}$, which may depend on treatments and covariates. For any model $({\bm{\delta}},{\bm{\gamma}})$, let the Kullback-Leibler optimal parameter value be
$\bm{\zeta}_{\delta,\gamma}^{*}=\arg\max_{\bm{\zeta}_{\delta,\gamma}}{\mathbb{E}}_{F^{*}}\left[\log p(y_{1}\mid\bm{\zeta}_{\delta,\gamma},\delta,\gamma)\right],$
and $H^{*}_{\delta,\gamma}$ its hessian evaluated at $\bm{\zeta}_{\delta,\gamma}=\bm{\zeta}_{\delta,\gamma}^{*}$.
14.1. Technical conditions
Let $\widehat{\bm{\zeta}}_{\delta,\gamma}$ be the MLE under model $({\bm{\delta}},{\bm{\gamma}})$ and
$A_{n}({\bf s})=\log p({\mathbf{y}}\mid\widehat{\bm{\zeta}}_{\delta,\gamma},{\bm{\delta}},{\bm{\gamma}})-\log p({\mathbf{y}}\mid\widehat{\bm{\zeta}}_{\delta,\gamma}+{\bf s}/\sqrt{n},{\bm{\delta}},{\bm{\gamma}})=-\frac{{\bf s}^{\top}H(\bm{\zeta}^{*}_{\delta,\gamma}){\bf s}}{2n}+r_{n}({\bf s}/\sqrt{n}),$
where $r_{n}$ is the error in the second-order Taylor log-likelihood expansion at $\bm{\zeta}_{\delta,\gamma}^{*}$. Let $H(\bm{\zeta}_{\delta,\gamma})$ be the log-likelihood hessian evaluated at $\bm{\zeta}_{\delta,\gamma}$. We assume that
  1. D0.
    $P\left(A_{n}({\bf s})\text{ is convex in }{\bf s}\right)\to 1$, as $n\to\infty$
  2. D1.
    $\widehat{\bm{\zeta}}_{\delta,\gamma}\stackrel{{\scriptstyle P}}{{\longrightarrow}}\bm{\zeta}_{\delta,\gamma}^{*}$ under $F^{*}$, as $n\rightarrow\infty$.
  3. D2.
    $H(\bm{\zeta}_{\delta,\gamma}^{*})/n\stackrel{{\scriptstyle P}}{{\longrightarrow}}H_{\delta,\gamma}^{*}$ under $F^{*}$, as $n\rightarrow\infty$, for a strictly positive-definite $H^{*}_{\delta,\gamma}$.
  4. D3.
    $\min_{\alpha_{t}^{*}\neq 0}|\alpha_{t}^{*}|\geq a$ and $\min_{\gamma_{j}^{*}\neq 0}|\beta_{j}^{*}|\geq b$ for some constants $a,b>0$.
Condition D0 requires that the log-likelihood is convex around the MLE, which holds with probability 1 at any ${\bm{\zeta}}_{\delta,\gamma}$ for generalized linear models with the canonical link. Conditions D1-D2 are minimal. If one assumes that $\bm{\zeta}_{\delta,\gamma}$ has bounded support, then D1 holds (Hjort and Pollard, 2011) and D2 also holds provided $|H(\bm{\zeta}_{\delta,\gamma}^{*})|$ has finite mean, by the continuous mapping theorem and strong law of large numbers. More generally one may show the asymptotic validity of a Taylor log-likelihood expansion around $\bm{\zeta}_{\delta,\gamma}^{*}$ and establish asymptotic normality of $\widehat{\bm{\zeta}}_{\delta,\gamma}$, see Theorem 4.1 in Hjort and Pollard (2011). Condition D3 is a beta-min condition that can be relaxed to allow for vanishing $(a,b)$, as long as $(a,b)$ are larger than $\sqrt{n}^{-1/2}$ times logarithmic terms, but here we assume fixed $(a,b)$ to obtain simpler expressions for the asymptotic Bayes factor rates.
14.2. Laplace approximation to Bayes factors
Assuming Conditions D0-D2, plus a prior boundedness condition that is satisfied in our setting, Proposition S6 in Rossell and Rubio (2021) gives that
(27) $\frac{p^{N}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})/p^{N}({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}{[\hat{p}^{N}({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}},{\bm{\theta}})/\hat{p}^{N}({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*},{\bm{\theta}})]\frac{p({\bm{\delta}})p({\bm{\gamma}}\mid{\bm{\theta}})}{p({\bm{\delta}}^{*})p({\bm{\gamma}}^{*}\mid{\bm{\theta}})}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}1$
as $n\rightarrow\infty$, where
(28) $\frac{\hat{p}^{N}({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}},{\bm{\theta}})}{\hat{p}^{N}({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*},{\bm{\theta}})}=e^{L({\bm{\delta}},{\bm{\gamma}})/2}\frac{p(\bm{\zeta}_{\gamma}^{*}\mid{\bm{\gamma}},{\bm{\delta}})}{p(\bm{\zeta}_{\gamma^{*}}^{*}\mid{\bm{\gamma}}^{*},{\bm{\delta}}^{*})}\left(\frac{2\pi}{n}\right)^{d/2}\frac{|H^{*}_{\delta^{*},\gamma^{*}}|^{1/2}}{|H^{*}_{\delta,\gamma}|^{1/2}},$
is a Laplace approximation to the Bayes factor between $({\bm{\delta}},{\bm{\gamma}})$ and $({\bm{\delta}}^{*},{\bm{\gamma}}^{*})$ under the Normal prior, $L({\bm{\delta}},{\bm{\gamma}})$ is the corresponding likelihood-ratio test statistic, and $d$ is the difference between the number of non-zero parameters in $({\bm{\delta}},{\bm{\gamma}})$ and $({\bm{\delta}}^{*},{\bm{\gamma}}^{*})$.
Therefore, in our asymptotic study we may replace the Bayes factor under Normal priors on the right-hand side of (26) by its Laplace approximation in 28.
14.3. Bayes factor rates
The frequentist properties of the ratio of posterior expectations in (26) and the Laplace approximation to the Bayes factor in (28) have been well-studied, e.g. see Rossell and Rubio (2018) (Proposition 5) for Gaussian outcomes and Rossell and Rubio (2021) (Propositions 3-4) for certain survival and generalized linear models. We now summarize the results. When $({\bm{\delta}},{\bm{\gamma}})$ is an overfitted model, combining Expressions (26) and (27), one may show that $p({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}})/p({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*})=(n\tau)^{-3d/2}\times O_{p}(1)$, where $d=|{\bm{\delta}}|_{0}+|{\bm{\gamma}}|_{0}-|{\bm{\delta}}^{*}|_{0}-|{\bm{\gamma}}^{*}|_{0}$ is the difference between model dimensions. In contrast, when $({\bm{\delta}},{\bm{\gamma}})$ is a non-overfitted model, then
(29) $\log\frac{p({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}})}{p({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*})}=-\frac{3d}{2}\log(n\tau)-nc+O_{p}(1),$
where $c>0$ is a constant that depends on $({\bm{\alpha}}^{*},{\bm{\beta}}^{*})$. Under Condition (D3), $c$ can be taken to be a fixed constant (i.e. not depending on $n$).
These Bayes factor rates, combined with the bounds on $\pi_{j}({\bm{\theta}})\in[\underline{\rho},\bar{\rho}]$, give Bayes factor rates for CIL under any given hyper-parameter ${\bm{\theta}}$. Below, we denote by $d_{1}=\sum_{j=1}^{J}\gamma_{j}(1-\gamma_{j}^{*})$ the number of covariates included in ${\bm{\gamma}}$ but not in ${\bm{\gamma}}^{*}$, and by $d_{2}=\sum_{j=1}^{J}(1-\gamma_{j})\gamma_{j}^{*}$ that of covariates included in ${\bm{\gamma}}^{*}$ but not in ${\bm{\gamma}}$.
Consider first overfitted models. Using that
$\frac{p({\mathbf{y}}\mid{\bm{\delta}},{\bm{\gamma}})}{p({\mathbf{y}}\mid{\bm{\delta}}^{*},{\bm{\gamma}}^{*})}=(n\tau)^{-3d/2}\times O_{p}(1),$
and that $p({\bm{\delta}})=p({\bm{\delta}}^{*})$ under our prior, we obtain
$\frac{p({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})}{p({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}$ $=(n\tau)^{-3d/2}\prod_{\gamma_{j}=1,\gamma_{j}^{*}=0}\frac{\pi_{j}({\bm{\theta}})}{1-\pi_{j}({\bm{\theta}})}\prod_{\gamma_{j}=0,\gamma_{j}^{*}=1}\frac{1-\pi_{j}({\bm{\theta}})}{\pi_{j}({\bm{\theta}})}\times O_{p}(1)$
$\leq(n\tau)^{-3d/2}\left(\frac{\bar{\rho}}{1-\bar{\rho}}\right)^{d_{1}}\left(\frac{1-\underline{\rho}}{\underline{\rho}}\right)^{d_{2}}\times O_{p}(1).$
Note that for over-fitted models $d_{1}=d$ and $d_{2}=0$, giving
$\frac{p({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})}{p({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}\leq(n\tau)^{-3d/2}\left(\frac{\bar{\rho}}{1-\bar{\rho}}\right)^{d}\times O_{p}(1),$
as we wished to show.
Consider now non-overfitted models. Using (29) and that $p({\bm{\delta}})=p({\bm{\delta}}^{*})$ gives
$\log\left(\frac{p({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})}{p({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}\right)=-\frac{3d}{2}\log(n\tau)-nc$
$+\log\left(\prod_{\gamma_{j}=1,\gamma_{j}^{*}=0}\frac{\pi_{j}({\bm{\theta}})}{1-\pi_{j}({\bm{\theta}})}\right)+\log\left(\prod_{\gamma_{j}=0,\gamma_{j}^{*}=1}\frac{1-\pi_{j}({\bm{\theta}})}{\pi_{j}({\bm{\theta}})}\right)+O_{p}(1).$
Noting that $\pi_{j}({\bm{\theta}})\in[\underline{\rho},\bar{\rho}]$, that there are $d_{1}$ terms such that $(\gamma_{j}=1,\gamma_{j}^{*}=0)$, and that there are $d_{2}$ terms such that $(\gamma_{j}=0,\gamma_{j}^{*}=1)$, gives
$\log\left(\frac{p({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})}{p({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}\right)\leq-\frac{3d}{2}\log(n\tau)-nc+d_{1}\log\left(\frac{\bar{\rho}}{1-\bar{\rho}}\right)+d_{2}\log\left(\frac{1-\underline{\rho}}{\underline{\rho}}\right)+O_{p}(1)$
as we wished to prove.
15. Supplementary Results
15.1. Illustration of the EB and EP objective functions